Can you please help me with this problem?
Solve the equation by isolating the natural logarithm and exponentiating both sides. Express the answer in terms of e.
6 + 6 ln x = 11

Thank You!!

Answer by stanbon(75887)

6 + 6 ln x = 11
6lnx=5
lnx=5/6
x=e^(5/6)
x=2.30097589...
Cheers,
Stan H.